write each equation in slope-intercept form
y+3 = -3(x -6)
thank you.

Answer by stanbon(75887)

write each equation in slope-intercept form
y+3 = -3(x -6)
---
y+3 = -3x+18
y = -3x+15
That is the slope-intercept form.
==================
Cheers,
Stan H.